# Large Knowledge Model: Perspectives and Challenges

### v1

@@ 2.4 Knowledge Editing
Knowledge editing EasyEdit is geared towards the precise modification of parameterized knowledge of large models. This process aims to rectify incorrect knowledge, eliminate harmful content, or update outdated information. Yet, executing edits in the vast parameter space of large-scale neural models is a nontrivial task. It necessitates firstly locating the incorrect knowledge, then secondly determining the scopes and boundaries of necessary modifications. Critically, because the parameters of the neural network are highly correlated, changing any one of the neurons will have a “pull-start” effect on the other neurons associated with it. The relationship between facts explicitly defined in a symbolic knowledge graph can be used to help determine the correlation relationship between parametric knowledge and define the boundaries of parameter updates. This understanding assists in delineating relationships for parameter updates, ensuring that modifications are both effective and do not inadvertently disrupt related knowledge structures EditLLMs .
2.5 Knowledge Alignment
Broadly speaking, since the majority of information within most of KGs has undergone manual review and verification, thereby ensuring a high degree of standardization and accuracy, it can serve as a target for aligning large models zhang2023knowledgeable . For instance, KGs can be instrumental in generating feedback signals for the reward model . This is achieved by contrasting the outputs of the LLM with the factual data, structured relations, or logic constraints present in the KG. Such a comparison not only aids in enhancing the accuracy and reliability of the outputs generated by the LLM but also provides valuable metrics to assess the model’s performance.
@@ 2.6 Semantic Tools for LLM
Tool-Augmented Language Models ALMSurvey ; Toolformer are proposed to extend the functionalities of large models by integrating the ability to invoke external tools for problems solving beyond their native solution scope. These external tools can range from mathematical calculators, database interfaces to specific functional APIs. A complex task may require a orchestration of calling multiple tools, forming a complex tool calling logic. For LLMs to adeptly master such complex tool-calling logic remains a significant challenge. Knowledge Graphs can play a valuable role in this context. They can enhance the semantic descriptions or model the calling relationships of these tools, thereby facilitating the modeling of complex API combination logic. This approach has already seen considerable application in the realm of traditional semantic web services mcilraith2002adapting , suggesting a promising avenue for enriching the functional versatility of tool-augmented large models.
@@ 2.7 Knowledgeable AI Agents
3 LLM-enhanced Knowledge Bases
@@ 3.1 LLM as KG Builder
In critical scenarios requiring high reliability, the construction of a symbol-based knowledge base would ensure accuracy, offer traceable sources for its content, and provide enhanced interpretability. LLMs with their proficiency in language understanding, AI-Generated content, and generalization capabilities, marks a new era in developing symbolic knowledge bases and their utility in serious or critical applications.
@@
  • Enhanced Generalization in Knowledge Extraction: As highlighted earlier, LLMs demonstrate remarkable precision in identifying previously unseen entity types and relations. This high level of accuracy is largely attributed to the models’ advanced generalization capabilities. This adaptability enables LLMs to effectively process and recognize novel data inputs, transcending the limitations of traditional extraction methods heavily relying on laborious labeling efforts.
  • AI-Generated Content for Knowledge Graph Construction: LLMs are equipped with powerful generative capabilities. In cases where knowledge is absent or insufficient in the model or source text, large models can still synthesize new knowledge for the KG. However, it is notable that while AIGC extends the depth and breadth of KGs, it also carries the risk of introducing incorrect knowledge.
@@ 3.2 LLM as KG Controller
  • Logical Query Parsing: Leveraging the language comprehension abilities of large models to convert natural language query into structured query formats (like SPARQL, Cypher, Gremlin, etc.) offers a more direct solution CoK ; NLQxform . This method frees programmers from the tedious and laborious task of manually devising query compositions, streamlining the process significantly.
  • Embedding-based Retrieval Augmentation EmbedKGQA : Lastly, large models can enhance the retrieval process from structured knowledge bases by utilizing embedding-based techniques. This involves representing queries and the knowledge base in a high-dimensional space to improve the accuracy and relevance of the retrieved information.
@@
3.3 Structured Knowledge Pretraining and Large Graph Model
Research shows that current large models often do not outperform traditional small models in tasks inherently reliant on structured knowledge, such as link prediction, association graph mining, subgraph analysis, or time-series prediction LLMsForKGC ; Evaluation_on_Relations . The Large Graph Models GraphFundationModel ; GraphGPT are then proposed for deriving insights from complex, graph-structured data, which is critical in a variety of advanced applications, including social network analysis, biological network analysis, and network security, among others.
In this vein, we can develop special pre-training mechanisms tailored for encoding structured knowledge, transforming structured knowledge into instruction formats to boost model generalizability, or devising specific reward models for graph computation tasks. This approach is particularly valuable for big data analysis applications dependent on graph-structured data.
KGTransformer KGTransformer exemplifies this approach. It employs several basic tasks for a KG, such as triple classification or head/tail prediction, to pretrain a structured knowledge graph. For downstream applications such as image classification, KGTransformer integrates a prompt fine-tuning mechanisms, which effectively harnesses the knowledge from the pretrained KG, thus enhancing downstream tasks’ performance.
3.4 LLM-enhanced Symbolic Reasoning
@@
“Reasoning” refers to the cognitive process in which people derive inferences or conclusions based on observed phenomena or facts, combined with known facts, experiences, or rules ReasoningWithLLMs ; ENeSy . Reasoning is one the most fundamental challenge in AI and has been extensively studied in many fields such as philosophy, computer science, and psychology. Due to the broad concept of reasoning, this article only discusses reasoning within the scope of language models and knowledge graphs.
@@
KGs and LLMs are both capable of performing reasoning. LLMs support a variety of reasoning tasks, such as mathematical reasoning, commonsense reasoning, symbolic manipulation, logical reasoning, and multi-modal reasoning ReasoningWithLLMs . In the realm of KGs, typical reasoning forms encompass concept reasoning, knowledge completion, multi-hop query reasoning, rule learning and rule-based reasoning, inconsistency reasoning, and analogical reasoning KGReasonWithLogicEmb ; AnalogyReason ; Ruleformer ; ENeSy .
The reasoning supported by KGs relies mainly on explicitly acquired symbolic knowledge, hence it is more interpretable, and the reasoning process is relatively reliable. With the integration of methods such as graph neural networks or KG embedding, reasoning can also be carried out in a vector space. However, since the knowledge coverage of a KG is not as extensive as that of LLMs, the robustness and generalization ability are relatively poor.
The reasoning supported by LLM ReasoningWithLLMs is mainly performed in a parameterized space, hence it is more of an implicit form of reasoning. Different from traditional reasoning methods such as graph neural network or KG embedding, LLM also rely on the richness of knowledge embedded in the large models. Furthermore, many studies show that the reasoning capabilities of large models are strongly related to the model size EmergentOfLLMs ; GPT3 . Certain reasoning abilities can only be activated when the model size reaches a certain scale. Since large models possess very good generalization abilities, they are particularly adept at extrapolating to new reasoning tasks. However, since the reasoning relies on a generative process, it is prone to hallucination issues, leading to unreliability in the reasoning results.
@@
Therefore, LLMs and KGs can complement with each other to achieve a more robust reasoning solution. On the one hand, a large language model can use a symbolic knowledge graph as an externally accessible tool. When it is necessary to give a definitive, reliable answer, the reasoning with a symbolic knowledge graph can be employed. On the other hand, when symbolic knowledge graph reasoning cannot derive an answer due to lack of knowledge, it is possible to use the extensive knowledge and high generalization ability of the large language model to give coarse-grained reasoning. Additionally, the interplay between LLMs (parametric, yet not entirely reliable) and KGs (symbolic, more interpretable) allows for their mutual enhancement and complementation.

### v2

@@ 2.6 Large Model Knowledge Editing
Knowledge editing EasyEdit is geared towards this purpose and aims to rectify incorrect knowledge, eliminate harmful content, or update outdated information parametrically. Yet, executing edits in the vast parameter space of large-scale neural models is a nontrivial task. It necessitates firstly locating the incorrect knowledge, then determining the scopes and boundaries of necessary modifications. Critically, because the knowledge within these models is highly interconnected, changing one piece of information could have far-reaching effects.
Efforts have been made to leverage knowledge graphs for generating structured data and enhancing reward feedback signals for large model optimization. The KnowPAT KnowPAT framework introduces a novel approach by integrating domain-specific knowledge graphs into the model alignment process, enabling knowledge-informed human-machine alignment. This method unfolds in three succinct steps: first, it retrieves relevant information from the knowledge graph based on the input query; second, it constructs a knowledge preference set by using knowledge triples of varied quality to align model outputs with human preferences; finally, it applies this preference set to fine-tune and align the model, improving its performance and alignment with human knowledge and preferences.
Figure 7: Editing knowledge in a LLM and a KG is quite different detox ; studyofedit
Knowledge graphs could enhance large model knowledge editing in three ways. First, the validated, accurate content serves as a prime reference input for edits. Second, they help navigate the complex logical relationships essential for modifying model knowledge, offering a logic roadmap for adjusting parameters. However, unlike knowledge graphs, changes in large models’ knowledge don’t always revert cleanly due to their attention-based encoding, leading to potential errors or hallucinations. The most effective, albeit challenging, method involves using knowledge graphs from the start—during pre-training or fine-tuning—to guide the organization and storage of model knowledge, closely mirroring the symbolic representation of knowledge graphs. This would not only improve knowledge connectivity within the model but also simplifies the editing process by clearly defining the scope and effects of modifications EditLLMs .
2.6.1 Knowledge-informed Human Alignment
Aligning large models with human expectations, a strategy to counter hallucinations and misinformation, involves tailoring model outputs to align with human values and ethics. This typically starts with supervised fine-tuning using curated responses, followed by ranking responses to train a reward model that reflects human value priorities. This reward model then informs a reinforcement learning process, optimizing the model’s alignment with human values—a crucial step toward aligning with human moral knowledge.
Knowledge Graphs (KGs), renowned for their high standardization and accuracy due to thorough manual review, offer a solid foundation for aligning large models. This high degree of reliability makes KGs ideal for generating feedback signals to inform the reward model, by comparing Large Language Model (LLM) outputs against KG’s factual data, structured relationships, and logical constraints. This comparison not only boosts the LLM’s output accuracy and reliability but also provides metrics to evaluate model performance.
Figure 8: KnowPAT: an example of improving human alignment with a KG KnowPAT
Building on this, as shown in Figure 8, the KnowPAT KnowPAT framework leverages KGs to optimize large model alignment through the creation of partial order training data and enhanced reward feedback signals. KnowPAT’s innovative strategy integrates domain-specific KGs into the alignment process for knowledge-driven human alignment. This process unfolds in three steps: firstly, extracting relevant KG information based on input queries; secondly, constructing a knowledge preference set from varied quality knowledge triples to align LLM outputs with human preferences; and finally, using this preference set for model fine-tuning, thereby enhancing both performance and alignment with human knowledge and preferences.
OpenAI’s Weak-to-Strong generalization weak-to-strong reveals a future where AI surpasses human ability, highlighting a new challenge of guiding more powerful AIs with weaker human input. This contrasts with traditional models where a stronger human teacher supervises a weak AI. Knowledge graphs, while limited in scope, offer a targeted approach to align AI with human knowledge. The challenge lies in using these precise but smaller, weaker knowledge graphs to guide broader, less accurate AI models. This research direction focuses on refining AI understanding and application of human knowledge through detailed knowledge graphs, striving for AI systems that are both powerful and aligned with human values.
@@ 2.7 Semantic Tools for LLM
Tool-Augmented Language Models ALMSurvey ; Toolformer are proposed to extend the functionalities of large models by integrating the ability to invoke external tools for problems solving beyond their native solution scope. These external tools can range from mathematical calculators, database interfaces to specific functional APIs. A complex task may require a orchestration of calling multiple tools, forming a complex tool calling logic.
Large models still face difficulties in autonomously learning such complex tool invocation logic, whereas knowledge graphs can be used in modeling the logic behind complex API combinations, a practice already prevalent in the traditional service computing architecture field mcilraith2002adapting . However, compared to the traditional software engineering field, by leveraging knowledge to model the logic of API calls and further guiding the process of API combination, with the language understanding capabilities of large models, we might not need to model the API relationships in very fine details. We might just need to provide textual descriptions of the tool’s functions and how to invoke them, along with a simple description of the relationships between tools, such as the tool’s hierarchical classification and simple associations. This could allow large models to learn the reasonable logic of tool invocation and combination on their own.
@@ 2.8 Knowledgeable AI Agents
Figure 9: KnowAgent: an example of knowledgeable AI agent knowagent
@@
2.9 KG-Augmented Large Multi-Modal Models
Large models are advancing towards integrating multiple modalities, enabling AI to master skills like seeing, hearing, and speaking, and enhancing interactions with the physical world. There has been substantial research on enhancing multi-modal processing with KGs. For instance, in zero-shot visual reasoning tasks, knowledge graphs are widely used to establish connections between new types of entities and known ones, assisting models to quickly gain the ability to recognize new types of entities by using knowledge graphs as a bridge for transfer learning knowledge-zero-shot .
Knowledge graphs can certainly be used to enhance the multi-modal generative capabilities of large models.For example, we can retrieve triples related to prompts from an external knowledge graph, and then use a large model to expand on these triples based on user input, enabling the large model to generate images that are richer in content and more in line with human expectations. In fact, the existing multi-modal large models are still quite weak in generating content that aligns with human cognitive concepts and intentions. We believe one reason for this is the lack of knowledge to constrain and control the generation process in these multi-modal large models. It’s similar to how a human painter first forms a conceptual understanding of the objective world, and then uses the logical relationships between these concepts to constrain the creative process, ensuring that the content produced aligns with human cognition.
Figure 10: Knowledge Graphs Meet Multi-modal Learning kgmm
3 LLM-enhanced Symbolic Knowledge Bases
@@
3.1 Symbolic Knowledge Bases: A Necessity
Large language models like GPT showcase impressive knowledge representation and processing skills. However, they are not a one-size-fits-all solution. In many practical settings, symbolic systems prove to be sufficient or even superior. In critical scenarios requiring high reliability, the construction of a symbol-based knowledge base would ensure accuracy, offer traceable sources for its content, and provide enhanced interpretability.
Importantly, most existing information systems are developed using symbolic representations. Often, there is no need to convert these into neural network representations for integration into large models, which can lead to a loss of transparency due to their “black box” nature. Looking forward, the future of intelligent systems likely lies in a hybrid approach that combines neural networks and symbolic systems, leveraging the strengths of each. A practical example of this is enhancing traditional symbolic knowledge bases with large models, which could be a more viable and scalable application in many real-world scenarios.
@@ 3.2 LLM as KG Builder
  • Instruction-driven Knowledge Extraction: As highlighted earlier, LLMs demonstrate remarkable precision in identifying previously unseen entity types and relations. This high level of accuracy is largely attributed to the advanced generalization capabilities enabled by instruction-driven learning. This adaptability enables LLMs to effectively process and recognize novel data inputs, transcending the limitations of traditional extraction methods heavily relying on laborious labeling efforts.
  • Generative Knowledge Graph Constructions: LLMs are equipped with powerful generative capabilities. In cases where knowledge is absent or insufficient in the model or source text, large models can still synthesize new knowledge for the KG. However, it is notable that while AIGC extends the depth and breadth of KGs, it also carries the risk of introducing incorrect knowledge.
@@ 3.3 LLM as KG Controller
  • Text to Logical Query Parsing: Leveraging the language comprehension abilities of large models to convert natural language query into structured query formats (like SPARQL, Cypher, Gremlin, etc.) offers a more direct solution CoK ; NLQxform . This method frees programmers from the tedious and laborious task of manually devising query compositions, streamlining the process significantly.
  • RAG-based KBQA EmbedKGQA : Lastly, large models can enhance the retrieval process from structured knowledge bases by utilizing embedding-based techniques. This involves representing queries and the knowledge base in a high-dimensional space to improve the accuracy and relevance of the retrieved information.
@@
3.4 Boosting Symbolic Reasoning with Large Models
3.4.1 LLM Reasoning vs KG Reasoning
@@
“Reasoning” refers to the cognitive process by which people derive conclusions from observed phenomena, combined with known facts, experiences, or rules ReasoningWithLLMs ; ENeSy . This fundamental AI challenge is widely studied across philosophy, computer science, and psychology. This article, however, focuses solely on reasoning within language models and knowledge graphs.
@@
Both Knowledge Graphs (KGs) and Large Language Models (LLMs) support varied reasoning tasks. LLMs handle mathematical, commonsense, symbolic, logical, and multi-modal reasoning ReasoningWithLLMs . KGs engage in concept reasoning, knowledge completion, multi-hop query reasoning, rule learning, inconsistency reasoning, and analogical reasoning KGReasonWithLogicEmb ; AnalogyReason ; Ruleformer ; ENeSy .
Reasoning in KGs is based on explicitly acquired symbolic knowledge, making it interpretable and reliable. Techniques like graph neural networks or KG embedding enable reasoning in a vector space, yet KGs often lack the robustness and generalization capabilities of LLMs due to limited knowledge coverage. LLM reasoning ReasoningWithLLMs occurs in a parameterized, implicit space, depending heavily on the embedded knowledge. Unlike traditional methods, the reasoning abilities of LLMs scale with model size EmergentOfLLMs ; GPT3 , enhancing their effectiveness on new tasks. However, this generative reasoning can lead to hallucinations, undermining result reliability.
LLM reasoning primarily focuses on leveraging textual semantics, while KG reasoning incorporates both textual semantics and the graph’s structural knowledge. Here, we particularly explore how large models can enhance KG reasoning, which can be broadly categorized into three strategies as elaborated in the following sections
3.4.2 LLM-Centric Reasoning with KG
The first is LLM-Centric Reasoning which emphasizes the inference capabilities of large models to enhance KG reasoning directly. A simple approach in this vein is to reason with knowledge graphs with prompt engineering, where structured data from the knowledge graph is presented to a large model as contextual prompts for in-context learning. However, studies have shown that this method often performs even worse than traditional models like TransE, especially when textual information about entities is not included LLMsForKGC ; KICGPT . This underperformance underscores the current large models’ limited ability to effectively learn from structured data, leading to the development of new models that are specifically trained on graph data for improved performance.
Prompt engineering involves creating complex prompts, but large language models (LLMs) might struggle to grasp the structured data in knowledge graphs. A better strategy is to use the structured data from knowledge graphs to create instructions and employ instruction fine-tuning to improve LLM training KnowPAT . An example is illustrated in Figure 11. This approach integrates structured signals with instructional data to enhance the model’s understanding of graph structures. It starts with pre-training using traditional KG embedding models to obtain initial entity and relationship embeddings, which include structured information. These embeddings are then combined with textual instructions for fine-tuning the LLM. Models fine-tuned this way typically outperform those without structural tuning and can even surpass traditional KG embedding models in tasks sensitive to structural details. The improved performance stems from the LLMs’ strong language understanding abilities and their capacity to process structured information, enriched further by the embedded structured knowledge in the models.
Figure 11: An example of augmenting LLM reasoning with a KG KoPA
3.4.3 KG-Centric Reasoning with LLM
The second is KG-Centric Reasoning which utilizes large models’ language understanding abilities and structured training mechanisms to improve the reasoning performed by the knowledge graphs themselves.
This approach builds on traditional knowledge graph reasoning models such as KG Embedding, graph neural networks, and classical symbolic reasoning. However, it enhances these models by integrating the capabilities of large language models, which serve multiple supportive functions. Firstly, large models greatly enhance the understanding of language and semantics within the knowledge graph, improving how textual semantics are processed. Additionally, they enrich the reasoning process by integrating commonsense knowledge inherent in their training data. Furthermore, large models contribute by generating diverse and plausible reasoning chains, adding depth and breadth to the inference paths available within the knowledge graph. This integration of large models enhances the traditional reasoning capabilities, making the processes more dynamic and contextually rich.
Another approach in this vein is structured knowledge pre-training, which we will explore further in upcoming discussions on large graph models. This method does not strictly require the involvement of LLM. Instead, it utilizes architectures like Transformers to pre-train on knowledge graphs. The model, trained entirely on structured signals, then performs reasoning on the knowledge graphs. KGTransformer KGTransformer exemplifies this approach. It employs several basic tasks for a KG, such as triple classification or head/tail prediction, to pretrain a structured knowledge graph. For downstream applications such as image classification, KGTransformer integrates a prompt fine-tuning mechanisms, which effectively harnesses the knowledge from the pretrained KG, thus enhancing downstream tasks’ performance.
3.4.4 Synergistic Reasoning with KG and LLM
@@
Lastly, a collaborative strategy combines the strengths of both large models and knowledge graphs, allowing them to work together interactively to achieve more sophisticated reasoning. On the one hand, a large language model can use a symbolic knowledge graph as an externally accessible tool. When it is necessary to give a definitive, reliable answer, the reasoning with a symbolic knowledge graph can be employed. On the other hand, when symbolic knowledge graph reasoning cannot derive an answer due to lack of knowledge, it is possible to use the extensive knowledge and high generalization ability of the large language model to give coarse-grained reasoning. Additionally, the interplay between LLMs (parametric, yet not entirely reliable) and KGs (symbolic, more interpretable) allows for their mutual enhancement and complementation.
3.5 Large Graph Model and Structured Knowledge Pretraining
Research shows that current large models often do not outperform traditional small models in tasks inherently reliant on structured knowledge, such as link prediction, association graph mining, subgraph analysis, or time-series prediction LLMsForKGC ; Evaluation_on_Relations . The Large Graph Models GraphFundationModel ; GraphGPT are then proposed for deriving insights from complex, graph-structured data, which is critical in a variety of advanced applications, including social network analysis, biological network analysis, and network security, among others. In this vein, we can develop special pre-training mechanisms tailored for encoding structured knowledge, transforming structured knowledge into instruction formats to boost model generalizability, or devising specific reward models for graph computation tasks. This approach is particularly valuable for big data analysis applications dependent on graph-structured data.
The implementation of large-scale graph models can be divided into three main approaches as surveyed in details in this article GraphFundationModel .
  • Graph-Centric Method. The first approach employs Graph Neural Networks (GNN-based Model), focusing primarily on traditional GNN algorithms across two main phases: pre-training and fine-tuning. The core architecture often utilizes the traditional Message Passing model or a Transformer model specially designed for graph structures. The Transformer model leverages attention mechanisms to assess the interaction weights among elements within the graph. During the pre-training phase, methods such as contrastive loss can be used, where graph data is augmented, deformed, or modified to create contrastive learning samples. These self-enhanced contrastive samples are then used to pre-train the model. In the fine-tuning phase, a variety of downstream graph tasks can be addressed, and model parameters can be fine-tuned for these tasks using prompt-based instructions.
  • LLM-Centric Method. The second approach utilizes Large Language Models (LLM-based Model) and involves transforming graph data into sequences that mimic natural language for training with language models. This method encompasses two techniques: Graph-to-token and Graph-to-text. In the Graph-to-token technique, nodes and relationships from the graph structure are incorporated into natural language descriptions while retaining the token format of the graph. This allows the information to be processed by the language model while preserving the structural elements. Conversely, the Graph-to-text method translates graph descriptions into full-text narratives. This technique is particularly prevalent in knowledge graphs where each data triplet naturally corresponds to a sentence, facilitating direct conversion into text. Both methods convert inputs into natural language sequences, permitting the application of standard language model loss functions such as next-token prediction or masked prediction during pre-training. It’s critical to recognize that these methods may involve some semantic loss, as the rich logical and semantic details present in graph structures can diminish when reduced to sequential formats.
  • Synergistic Method. To mitigate the risk of semantic loss associated with relying solely on Large Language Models (LLMs), the third approach combines the capabilities of Graph Neural Networks (GNNs) and LLMs. In this hybrid model, LLMs can enhance GNNs by augmenting graph data before it is inputted into the GNN for prediction, leveraging the linguistic prowess of LLMs to enrich the data. Alternatively, GNNs can improve the performance of LLMs by providing structured signals that guide the reasoning process of the language models. This synergy can also extend to a bidirectional enhancement where both GNNs and LLMs align and refine each other’s predictions, improving overall accuracy and coherence in outcomes. This integrated approach harnesses the strengths of both systems to overcome their individual limitations and boost their predictive capabilities.
4 From LLM to LKM
Human knowledge, characterized by its complexity, is not optimally represented by mere sequences of free text. Our descriptions of the world often entail a variety of structured forms. One possible goal would be to create more advanced Large Knowledge Model (LKM) proficient in handling and deciphering the myriad structures of knowledge representation. This section aims to illuminate a roadmap and identify the key challenges in developing such a Large Knowledge Model.
4.1 Knowledge Soup: The Intricate Nature of Human Knowledge
The ability to process knowledge is one of the distinctive features that sets human intelligence apart from that of other species. For instance, lower life forms can only react directly to environmental stimuli, while smarter animals like cats and dogs are capable of selective actions based on perception, memory, and analogy. Humans, however, can abstract and induce knowledge from observations of the objective world and engage in complex reasoning through deduction, causation, analogy, and induction, thereby generating more rational behaviors. These higher-level intellectual activities are accomplished through knowledge, with natural language serving merely as one form of representing, carrying, and transmitting this knowledge. However, human knowledge is highly complex.
Figure 12: The Knowledge Soup Proposed by John Sowa knowledge-soup
In the early studies of artificial intelligence knowledge representation, a concept known as "Knowledge Soup" was introduced by the renowned knowledge engineering scholar John Sowa knowledge-soup . He believed that human knowledge plays a central role in cognitive behavior, but in reality, it is difficult to accurately represent and depict. He likened human knowledge to a bowl of soup: fluid rather than solidified, loosely organized rather than strictly defined, and dynamically changing rather than static. The "soup" contains large solid chunks, fragmented particles, and flowing, formless liquids. Due to external heating or human influence, large particles are broken down into smaller ones, which dissolve into the liquid, while the liquid itself continuously evaporates and changes. This is akin to having large chunks of multimodal knowledge as well as fine-grained triplet knowledge or structured, logical, conceptual, and rule-based knowledge, alongside the fluid and formless liquid that could represent parameterized neural network knowledge. We can fragment large blocks of text to create fine-grained knowledge and further integrate this knowledge with modern neural networks, blending various types of knowledge together as like a fluid manner.
4.2 Decoupling World Knowledge from Large Models
In fact, the RAG (Retrieval-Augmented Generation) approach can be seen as a method of separating the knowledge base from larger models. In this setup, domain-specific knowledge bases are maintained and trained independently of the larger models. This allows for more focused and specialized handling of knowledge within these specific areas.
4.3 Cognitive Alignment with Human Knowledge
Numerous cognitive scientists posit that humans inherently favor structured thinking, tending to recall, reason, and plan via associative thinking. Recent research suggests that the way large models store knowledge lacks discernible patterns as like a KG StructGPT . In other words, those knowledge that should be structured correlated does not show the expected associations in the parameter space, making it difficult for the model to make feasible correlations among knowledge pieces. There might be a sort of “cognitive gap” between the form of knowledge storage in large models and the organization of knowledge in the human mind. This cognitive gap becomes problematic in many critical scenarios when we expect the model to have more controllability, or to deal with deeper and more complex associative reasoning.
One way to alleviate this problem is to enhance the structure of the training corpus during the pretraining phase, known as restructure pretraining GraphGPT or structure-inducing pretraining StructIndPretrain . That is, we can use automated or semi-automated methods to enhance the structured features of the text. For example, we can organize sentences more logically (essentially, CoT is a method to increase the structure and logic of texts). We can structure paragraphs more coherently, add links between words, create connections between training samples, or directly incorporating an external structured KG to structurally reinforce the entire training corpus.
This raise up an interesting question: how essentially knowledge is stored and activated in large model? Research in this area, known as “circuit pathways” Circuit , aims to explore how the neural substructures are formed as parameters and incrementally activated during predictions. More specific studies seek to uncover how components of triplet facts are stored and activated FactRecall . Delving into the deep mechanisms of knowledge storage and reasoning in large language models is beyond this article’s scope, but it raises several important questions: what factors shape the knowledge circuits within these models? Is the knowledge in human brain organized similarly to text sequences? The answer is likely not, suggesting that learning and activating such knowledge would require more advanced training techniques.
We need to align the representation, storage, and reasoning mechanisms of large models with the human brain’s knowledge structures and cognitive processes. This alignment involves more than just adding structural signals or enhancing knowledge; it requires a fundamental overhaul of how knowledge is presented in training data, optimization functions, and constraints. By aligning the training data more closely with human cognitive structures, we can bridge the cognitive gap between large models and the human brain, enhancing the clarity, connectivity, and verifiability of the model’s knowledge. This transformation could effectively turn a Large Language Model (LLM) into a Large Knowledge Model (LKM), aligning its knowledge structure more naturally with human cognitive processes.